Information about direct materials cost follows for a local company:

Standard price per materials gram $ 16
Actual quantity used 2,250 grams
Standard quantity allowed for production 2,375 grams
Price variance $ 14,400 F

Required:

What was the actual purchase price per gram? (Round your answer to 2 decimal places.)
Answer #1

Price Variance = (SP-AP) *AQ

(16 - AP) *2250 = 14400

(16 - AP) = 14400/2250 = 6.40

AP = 16 - 6.40

= 9.60

Comment if you face any issues